If the percent chlorine in a salt crystal, KCl is 47.55 %, what is the percent chlorine in a kilogram of salt?
Answer #1

Mass of KCl = 74.55 g

Thus, 74.55 g of KCl has 47.55 % of Chlorine

Moles of KCl in 1kg = 1000 g of salt = mass/ Molar mass = 1000 g / (74.55 g/mol) = 13.41 mol

1 mol of KCl = 1 mol of Cl-

so, Moles of Cl- = 13.41 mol

Mass of Cl- = moles * molar mass = 13.41 mol* 35.453 g/mol = 475.56 g

Mass percent in a kilogram of salt = (Mass of Cl/ Molar mass of KCl)*100% = (475.56 g/ 1000 g)*100% = 47.55 %

Thus, Mass percent of chlorine remains the same .